40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 21 12:25:46 2019, gautam, Update, ALS, DFD electronics checkout ALSnoiseAnomaly.pdfDFDcalib.pdfspectra.pdfPTgains.pdf
    Reply  Tue Oct 22 15:32:15 2019, gautam, Update, ALS, EX uPDH electronics checkout uPDH_X_OLTFs.pdf
       Reply  Wed Oct 23 10:23:26 2019, rana, Update, ALS, EX uPDH electronics checkout 
          Reply  Wed Oct 23 11:11:01 2019, gautam, Update, ALS, EX uPDH electronics checkout 
    Reply  Fri Oct 25 01:04:49 2019, gautam, Update, ALS, ALS electronics chain was saturating 
       Reply  Mon Oct 28 23:20:11 2019, gautam, Update, ALS, ALS power budget 
Message ID: 14981     Entry time: Mon Oct 21 12:25:46 2019     Reply to this: 14984   14993
Author: gautam 
Type: Update 
Category: ALS 
Subject: DFD electronics checkout 


There are no unexpected red-flags in the performance of the DFD electronics. The calibration factors for the digital phase tracker system are 71.291 +/- 0.024 deg/MHz for the X delay line and 70.973 +/- 0.024 deg/MHz for the Y delay line, while the noise floor for the frequency noise discrimination is ~0.5 Hz/rtHz above 1 Hz (dominated by ADC noise).


  1. Attachment #1 - This observation is what motivated my investigation.
    •  found that for certain beat frequencies between the PSL + EX lasers, the frequency noise reported by the DFD system was surprisingly low.
    • The measurement condition was: EX laser frequency locked to the arm cavity length by the uPDH servo at EX, arm cavity length locked to PSL frequency via POX locking.
  2. To investigate further, I disconnected the output of the NF1611 PDs going to the ZHL-3A amplifiers on the PSL table (after first blocking the PSL light so that the PDs aren't generating any RF output).
    • An RF function generator (IFR2023B) was used to generate an RF signal to mimic the ALS beat signal.
    • I used a power splitter to divide the signal power equally between the two DFD paths.
    • The signal level on the Marconi was set to -5 dBm, to mimic the nominal power level seen by the DFD system.
    • I then performed two tests - (i) to calibrate the Phase Tracker output to deg / MHz and (ii) to measure the frequency noise reported by the DFD system for various signal frequencies.
    • Test (i): sweep the marconi frequency between 10 MHz - 200 MHz, measure the I and Q channels for each phase tracker servo, and figure out the complex argument of the signal using the arctangent. A linear polynomial was fit to the measured datapoints to extract the desired slope.
    • Test (ii): Sample frequencies uniformly distributed between 20 MHz - 80 MHz (nominal range of ALS beat frequencies expected). Reset the phase tracker servo gain, clear the output histories, wait for any transients to die out, and then collect the phase tracker servo output for 1 minute. Compute the FFT to figure out the frequency noise.
    • Attachment #2: Shows the phase tracker calibration, i.e. the results of Test (i). I took this opportunity to update the EPICS calibration fields that convert phase tracker servo output to Hz, the correction was ~7%. These numbers are consistent with what I measured previously - but the updated values weren't registered with SDF so everytime the LSC model was restarted, it reverted to the old values.
    • Attachment #3: Shows the spectra for the various measurements from Test (ii).
    • Attachment #4: Shows the gain of the phase tracker servo as a function of the RF signal frequency. This is a proxy for the signal strength, and the observed trend suggests that the signal power seen after digitization of the demodulated delay line output goes down by ~20% at 80 MHz relative to the level at 20 MHz. Seems reasonable to me, given frequency dependent losses of the intervening electronics / cabling.

Conclusion and next steps:

I still don't know what's responsible for the anomalously low noise levels reported by the ALS-X system sometimes. Next test is to check the EX PDH system, since on the evidence of these tests, the problem seems to be imprinted on the light (though I can't imagine how the noise becomes lower?).

Attachment 1: ALSnoiseAnomaly.pdf  76 kB  Uploaded Mon Oct 21 13:49:07 2019  | Hide | Hide all
Attachment 2: DFDcalib.pdf  165 kB  Uploaded Mon Oct 21 13:49:28 2019  | Hide | Hide all
Attachment 3: spectra.pdf  1.073 MB  Uploaded Mon Oct 21 15:30:57 2019  | Hide | Hide all
Attachment 4: PTgains.pdf  115 kB  Uploaded Mon Oct 21 15:31:11 2019  | Hide | Hide all
ELOG V3.1.3-